My Verbal Equation Is To Figure Out What:
Twelve less than 2 times a number is equal to -1 is.
Any help with this problem would be greatly appreciated=)
,Kia Jayy

Answer by rcmcc(152)

Start by making "a number"=x "twelve less"=-12 and "equal" to =
by changing the words into their literal meaning you get the equation
-12+2x=-1
add 12 to each side
-12+2x=-1
+12 +12
________________
2x=11
divide by 2
x=5.5 or 11/2
so your number is 5.5